OpenStudy (anonymous):

Which of the following will equal x^2+5x+6/x^2-x-6? When x does NOT equal -2 or 3. I got x+2/x-3. Am I correct?

OpenStudy (anonymous):

The choices are x+2/x-2 x+3/x-2 x+2/x-3 or x+3/x-3

OpenStudy (anonymous):

x+3/x-3

OpenStudy (anonymous):

all the condition means is that you can cancel factors

OpenStudy (anonymous):

(x+2)(x+3) / (x-3)(x+2)

OpenStudy (anonymous):

= (x+3)/(x-3)

OpenStudy (anonymous):

\[(x+3)/(x-3)\]

OpenStudy (anonymous):

s (x+3)/(x-3)

OpenStudy (anonymous):

see basically u hav to factorise the numerator and denominator, and then cancel out the common factors... numerator=(x+3)(x+2) denominator=(x-3)(x+2) so the common factor is (x+2), so calcel them and u get, (x+3)/(x-3) ps:see wen u hav 6/2 u write it as 3. how d'u do it? actually u express 6=3*2 so 3*2/2=3.u cancel the common factor 2 and get the answer. in the similar manner do wit these terms.